### v1

[…]
CHANDRASEKHAR GOKAVARAPU
[…]
This article develops a self-contained affine $\Gamma$-scheme theory for a class of commutative ternary $\Gamma$-semirings. By establishing all geometric and spectral results internally, the work provides a unified framework for triadic symmetry and spectral analysis. The central thesis is that a triadic $\Gamma$-algebra canonically induces two primary structures: (i) an intrinsic triadic symmetry in the sense of a Nambu–Filippov-type fundamental identity on the structure sheaf, and (ii) a canonical Laplacian on the finite $\Gamma$-spectrum whose spectral decomposition detects the clopen (connected-component) decomposition of the underlying space. We define $\Gamma$-ideals and prime $\Gamma$-ideals, endow $\operatorname{Spec_{\Gamma}}(T)$ with a $\Gamma$-Zariski topology, construct localizations and the structure sheaf on the basis of principal opens, and prove the affine anti-equivalence between commutative ternary $\Gamma$-semirings and affine $\Gamma$-schemes. Furthermore, we demonstrate that the triadic bracket on sections is invariant under $\Gamma$-automorphisms and compatible with localization. The main spectral theorem establishes the block-diagonalization of the Laplacian under topological decompositions and provides an algebraic-connectivity criterion. The theory is verified through explicit computations of finite $\Gamma$-spectra and their corresponding Laplacian spectra.
[…]
ternary $\Gamma$-semiring; $\Gamma$-spectrum; affine $\Gamma$-scheme; Nambu bracket; Laplacian; algebraic connectivity.
[…]
14A15; 16Y60; 05C50; 37J35.
1 Introduction:
The development is motivated by the need to bridge commutative ternary algebra with geometric symmetry The scope of this paper is as follows.
  • Symmetry / invariance. $\Gamma$-automorphisms of a ternary $\Gamma$-semiring act on the $\Gamma$-spectrum and on the structure sheaf. We exhibit an intrinsic triadic bracket on sections (a Nambu/Filippov-type structure) and prove its invariance under these actions [13, 14].
  • Geometry. We develop an affine $\Gamma$-scheme theory, including $\Gamma$-Zariski topology, localization, structure sheaf, and affine anti-equivalence, in the spirit of classical scheme theory [6, 7, 8, 9, 10] and relative/monoidal geometries [11, 12].
  • Spectral decomposition. From the (finite) $\Gamma$-spectrum we construct a canonical Laplacian based on specialization relations, prove invariance under $\Gamma$-scheme isomorphisms, and establish a block decomposition determined by clopen components. This yields an exact spectral factorization, with algebraic connectivity in the sense of Fiedler [15, 16].
Self-containedness. This is a standalone development. While related ideas exist for $\Gamma$-rings and ternary systems [1, 2, 3, 4, 5], all results required in the proofs below are provided here.
Roadmap. The development of the theory proceeds in four conceptual phases.
Phase I (Sections 2 and 3) establishes the affine $\Gamma$-scheme foundations: the $\Gamma$-spectrum, $\Gamma$-Zariski topology, standard-cover lemma, localization, structure sheaf, and the proof of the affine anti-equivalence.
Phase II (Section 4) introduces the triadic symmetry and Nambu-type structure on the structure sheaf and establishes its functoriality.
Phase III (Section 5) builds the bridge to spectral geometry by defining the canonical Laplacian and proving the block-diagonalization and algebraic-connectivity theorems.
Phase IV (Section 6) provides explicit computed examples to verify the theoretical results. Finally, Appendix A gives an optional fuzzy robustness viewpoint [20, 21, 22, 23, 24] and Appendix B sketches a conservative spectral clustering pipeline [18, 19, 17].
2 Phase I: affine $\Gamma$-scheme foundations
2.1 Commutative ternary $\Gamma$-semirings
We work with a concrete, scheme-friendly class of ternary $\Gamma$-semirings in which the $\Gamma$-parameter encodes a distinguished family of central units scaling a ternary multiplication. This choice yields (i) a nontrivial $\Gamma$-dependence, (ii) a clean localization theory, and (iii) a transparent geometric interpretation.
Definition 2.1 (Commutative semiring).
A commutative semiring is a tuple $(T,+,\cdot,0,1)$ where $(T,+,0)$ is a commutative monoid, $(T,\cdot,1)$ is a commutative monoid, multiplication distributes over addition, and 0 is absorbing: $0\cdot x=x\cdot 0=0$.
Definition 2.2 (Ternary $\Gamma$-semiring).
Let $(T,+,\cdot,0,1)$ be a commutative semiring and let $\Gamma$ be a commutative group. A (commutative) ternary $\Gamma$-semiring is the datum of a group homomorphism
$u:\Gamma\longrightarrow T^{\times},\qquad\gamma\longmapsto u_{\gamma},$
into the unit group of $T$, together with the induced $\Gamma$-parametrized ternary operation
$\{a\,b\,c\}_{\gamma}\ :=\ a\cdot b\cdot c\cdot u_{\gamma},\qquad a,b,c\in T,\ \gamma\in\Gamma.$
A homomorphism of ternary $\Gamma$-semirings $\varphi:(T,u)\to(S,v)$ is a semiring homomorphism $\varphi:T\to S$ such that $\varphi(u_{\gamma})=v_{\gamma}$ for all $\gamma\in\Gamma$.
Remark 2.3.
The ternary operation above is commutative in the three $T$-inputs and inherits associativity from the binary product. If $\Gamma$ is trivial, then $\{abc\}_{1}=abc$, recovering the usual ternary multiplication studied in ternary semiring theory (e.g. [4, 5]).
2.2 $\Gamma$-ideals, radicals, primes
Definition 2.4 ($\Gamma$-ideal).
Let $(T,u)$ be a ternary $\Gamma$-semiring. A subset $I\subseteq T$ is a $\Gamma$-ideal if:
  1. [label=(),leftmargin=2.2em]
  2. 1.
    $(I,+,0)$ is a submonoid of $(T,+,0)$, and
  3. 2.
    for all $a\in I$, all $b,c\in T$, and all $\gamma\in\Gamma$, one has $\{a\,b\,c\}_{\gamma}\in I$.
Equivalently, $I$ is an ideal of the underlying semiring $(T,+,\cdot,0,1)$ (since $u_{\gamma}$ is a unit for every $\gamma$).
Definition 2.5 (Generated $\Gamma$-ideal).
For $E\subseteq T$, the $\Gamma$-ideal generated by $E$, denoted $\langle E\rangle_{\Gamma}$, is the intersection of all $\Gamma$-ideals containing $E$. For a single element $f\in T$ we write $\langle f\rangle_{\Gamma}$.
Definition 2.6 ($\Gamma$-radical).
For a $\Gamma$-ideal $I$, define its $\Gamma$-radical by
$\sqrt[\Gamma]{\,\cdot\,}(I):=\{x\in T\mid\exists n\geq 1\text{ with }x^{n}\in I\},$
where $x^{n}$ is the $n$-fold binary product in the semiring.
Lemma 2.7.
For any $\Gamma$-ideal $I$, the set $\sqrt[\Gamma]{\,\cdot\,}(I)$ is a $\Gamma$-ideal containing $I$.
Proof 2.8.
If $x^{n}\in I$ then clearly $x\in\sqrt[\Gamma]{\,\cdot\,}(I)$; hence $I\subseteq\sqrt[\Gamma]{\,\cdot\,}(I)$. Closure under addition and multiplication by arbitrary elements is standard: if $x^{m}\in I$ and $y^{n}\in I$, then $(x+y)^{m+n}$ expands into a sum of monomials each divisible by $x^{m}$ or $y^{n}$, hence lies in $I$; similarly $(t\cdot x)^{m}=t^{m}x^{m}\in I$. Since $\Gamma$-ideal closure is equivalent to ideal closure in our setting (Definition 2.4), we are done.
Definition 2.9 (Prime $\Gamma$-ideal).
A proper $\Gamma$-ideal $P\subsetneq T$ is prime if for all $a,b,c\in T$ and $\gamma\in\Gamma$,
$\{a\,b\,c\}_{\gamma}\in P\quad\Longrightarrow\quad a\in P\ \text{or}\ b\in P\ \text{or}\ c\in P.$
Remark 2.10.
Because $u_{\gamma}$ is a unit, $\{abc\}_{\gamma}\in P$ is equivalent to $abc\in P$. Thus primeness here is equivalent to the binary statement:
$xy\in P\ \Rightarrow\ x\in P\ \text{or}\ y\in P,$
and hence coincides with the usual prime-ideal notion for commutative semirings. The ternary formulation is kept because it interacts transparently with the triadic bracket in Section 4.
Lemma 2.11 (Radical as intersection of primes).
Let $I$ be a $\Gamma$-ideal. Then
$\sqrt[\Gamma]{\,\cdot\,}(I)=\bigcap_{\begin{subarray}{c}P\ \text{prime }\Gamma\text{-ideal}\\ I\subseteq P\end{subarray}}P.$
Proof 2.12.
($\subseteq$) If $x\in\sqrt[\Gamma]{\,\cdot\,}(I)$, pick $n$ with $x^{n}\in I$. If $P$ is prime and contains $I$, then $x^{n}\in P$ implies $x\in P$. So $x$ lies in every prime containing $I$.
($\supseteq$) Suppose $x\notin\sqrt[\Gamma]{\,\cdot\,}(I)$. Consider the collection
$\mathcal{S}=\{J\mid J\text{ is a }\Gamma\text{-ideal},\ I\subseteq J,\ \text{and }x^{n}\notin J\ \forall n\geq 1\},$
partially ordered by inclusion. $\mathcal{S}$ is nonempty because $I\in\mathcal{S}$. By Zorn’s lemma, there exists a maximal element $P\in\mathcal{S}$.
We claim $P$ is prime. If $ab\in P$ and $a\notin P$, then the $\Gamma$-ideal $P+\langle a\rangle_{\Gamma}$ strictly contains $P$. By maximality, it must contain some $x^{m}$. Similarly, if $b\notin P$ then $P+\langle b\rangle_{\Gamma}$ contains some $x^{n}$. Multiplying the two containments gives $x^{m+n}\in P+\langle ab\rangle_{\Gamma}\subseteq P$ since $ab\in P$. This contradicts $P\in\mathcal{S}$. Hence $a\in P$ or $b\in P$, proving primeness.
Finally, $I\subseteq P$ by construction, and $x\notin P$ because $x\in P$ would force $x^{1}\in P$, contradicting $P\in\mathcal{S}$. Thus $x$ is not in the intersection of primes containing $I$.
2.3 The $\Gamma$-spectrum and $\Gamma$-Zariski topology
Definition 2.13 ($\Gamma$-spectrum).
The prime $\Gamma$-spectrum of $T$ is
$\operatorname{Spec_{\Gamma}}(T):=\{P\subsetneq T\mid P\ \text{is a prime }\Gamma\text{-ideal}\}.$
Definition 2.14 (Closed sets and principal opens).
For a $\Gamma$-ideal $I$, define
$V(I):=\{P\in\operatorname{Spec_{\Gamma}}(T)\mid I\subseteq P\}.$
For $f\in T$, define the principal open
$D_{\Gamma}(f):=\operatorname{Spec_{\Gamma}}(T)\setminus V(\langle f\rangle_{\Gamma})=\{P\in\operatorname{Spec_{\Gamma}}(T)\mid f\notin P\}.$
Proposition 2.15 ($\Gamma$-Zariski topology).
The assignment $I\mapsto V(I)$ satisfies:
$V(0)=\operatorname{Spec_{\Gamma}}(T),\quad V(T)=\emptyset,\quad V(I)\cup V(J)=V(I\cap J),\quad\bigcap_{\lambda}V(I_{\lambda})=V\Big(\sum_{\lambda}I_{\lambda}\Big).$
Hence the sets $V(I)$ define the closed sets of a topology on $\operatorname{Spec_{\Gamma}}(T)$, called the $\Gamma$-Zariski topology; the sets $D_{\Gamma}(f)$ form a basis of opens.
Proof 2.16.
Standard ideal-theoretic manipulations; the only point is that $\Gamma$-ideals form a complete lattice under intersection and sum, and the definition of $V(I)$ is inclusion-reversing. The verifications are identical to the ring case [7, 8].
Lemma 2.17 (Intersection of principal opens).
For $f,g\in T$ one has $D_{\Gamma}(f)\cap D_{\Gamma}(g)=D_{\Gamma}(fg)$.
Proof 2.18.
For a prime $\Gamma$-ideal $P$, $P\in D_{\Gamma}(f)\cap D_{\Gamma}(g)$ iff $f\notin P$ and $g\notin P$. Since $P$ is prime, this is equivalent to $fg\notin P$, i.e. $P\in D_{\Gamma}(fg)$.
2.4 The standard-cover equivalence lemma
The following is the key “cover vs. radical” equivalence required to run the sheaf/gluing theory on the basis of principal opens.
Theorem 2.19 (Standard-cover equivalence).
Let $f,f_{1},\dots,f_{n}\in T$. Then
$D_{\Gamma}(f)=\bigcup_{i=1}^{n}D_{\Gamma}(f_{i})\qquad\Longleftrightarrow\qquad f\in\sqrt[\Gamma]{\,\cdot\,}\big(\langle f_{1},\dots,f_{n}\rangle_{\Gamma}\big).$
Proof 2.20.
($\Rightarrow$) Let $I=\langle f_{1},\dots,f_{n}\rangle_{\Gamma}$. The equality of opens says: for every prime $P$, if $f\notin P$ then $f_{i}\notin P$ for some $i$. Equivalently, if $f_{i}\in P$ for all $i$ (i.e. $I\subseteq P$), then $f\in P$. Thus $f$ lies in the intersection of all primes containing $I$, hence $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ by Lemma 2.11.
($\Leftarrow$) Suppose $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$. Let $P\in D_{\Gamma}(f)$, so $f\notin P$. If $P$ contained $I$, then $f\in\sqrt[\Gamma]{\,\cdot\,}(I)\subseteq\sqrt[\Gamma]{\,\cdot\,}(P)=P$, contradicting $f\notin P$. Hence $I\not\subseteq P$, so some generator $f_{i}\notin P$, i.e. $P\in D_{\Gamma}(f_{i})$. Thus $D_{\Gamma}(f)\subseteq\bigcup_{i}D_{\Gamma}(f_{i})$, and the reverse inclusion holds because each $D_{\Gamma}(f_{i})\subseteq D_{\Gamma}(f)$ is automatic from the displayed radical inclusion: $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ implies $D_{\Gamma}(f)\supseteq D_{\Gamma}(f_{i})$ for each $i$.
Corollary 2.21.
If $D_{\Gamma}(f)=\bigcup_{i=1}^{n}D_{\Gamma}(f_{i})$, then there exists $N\geq 1$ and elements $a_{1},\dots,a_{n}\in T$ such that
$f^{N}=a_{1}f_{1}+\cdots+a_{n}f_{n}.$
Proof 2.22.
From Theorem 2.19 we have $f\in\sqrt[\Gamma]{\,\cdot\,}(\langle f_{1},\dots,f_{n}\rangle_{\Gamma})$, so $f^{N}\in\langle f_{1},\dots,f_{n}\rangle_{\Gamma}$ for some $N$. By definition of generated ideal, $f^{N}$ is a finite sum of multiples of the generators, giving the claim.
2.5 Localization and the structure presheaf
Definition 2.23 (Localization at a principal element).
Let $f\in T$ and let $S_{f}:=\{1,f,f^{2},\dots\}$. Define an equivalence relation on $T\times S_{f}$ by
$(a,s)\sim(b,t)\quad\Longleftrightarrow\quad\exists u\in S_{f}\text{ such that }u\cdot a\cdot t=u\cdot b\cdot s.$
Denote the equivalence class of $(a,s)$ by $\frac{a}{s}$, and set $T_{f}:=S_{f}^{-1}T:=(T\times S_{f})/\sim$. Addition and multiplication are defined by
$\frac{a}{s}+\frac{b}{t}:=\frac{at+bs}{st},\qquad\frac{a}{s}\cdot\frac{b}{t}:=\frac{ab}{st}.$
Lemma 2.24.
The operations in Definition 2.23 are well-defined and make $T_{f}$ a commutative semiring. The canonical map $\iota_{f}:T\to T_{f}$, $\iota_{f}(a)=\frac{a}{1}$, is a semiring homomorphism and $\iota_{f}(f)$ is a unit in $T_{f}$.
Proof 2.25.
Standard localization verifications for commutative semirings: the relation $\sim$ is compatible with the operations, and the formulas respect equivalence classes. The element $\iota_{f}(f)=\frac{f}{1}$ has inverse $\frac{1}{f}$ in $T_{f}$.
Proposition 2.26 (Universal property).
Let $\varphi:T\to R$ be a semiring homomorphism such that $\varphi(f)$ is a unit in $R$. Then there exists a unique semiring homomorphism $\widetilde{\varphi}:T_{f}\to R$ with $\widetilde{\varphi}\circ\iota_{f}=\varphi$, given by $\widetilde{\varphi}\big(\frac{a}{f^{n}}\big)=\varphi(a)\varphi(f)^{-n}$.
Proof 2.27.
Well-definedness follows from the defining relation $\sim$; uniqueness follows because every element of $T_{f}$ is represented by a fraction $\frac{a}{f^{n}}$.
Definition 2.28 (Structure presheaf on principal opens).
Define a presheaf on the basis $\{D_{\Gamma}(f)\}$ by
$\mathcal{O}_{T}\big(D_{\Gamma}(f)\big):=T_{f}.$
If $D_{\Gamma}(g)\subseteq D_{\Gamma}(f)$, then $f$ is invertible on $D_{\Gamma}(g)$, so by Proposition 2.26 there is a unique restriction homomorphism $\rho_{f,g}:T_{f}\to T_{g}$.
2.6 Sheaf condition on the basis
Theorem 2.29 (Sheaf condition on principal opens).
Let $f,f_{1},\dots,f_{n}\in T$ and assume $D_{\Gamma}(f)=\bigcup_{i=1}^{n}D_{\Gamma}(f_{i})$. Then the presheaf $\mathcal{O}_{T}$ satisfies:
  1. [label=(),leftmargin=2.2em]
  2. 1.
    (existence of gluing) For any sections $s_{i}\in T_{f_{i}}$ such that their restrictions agree on each overlap $D_{\Gamma}(f_{i})\cap D_{\Gamma}(f_{j})=D_{\Gamma}(f_{i}f_{j})$, there exists $s\in T_{f}$ with $s|_{D_{\Gamma}(f_{i})}=s_{i}$ for all $i$.
  3. 2.
    (uniqueness) Such an $s$ is unique.
Proof 2.30.
Step 1: choosing a common power relation. By Corollary 2.21 there exist $N\geq 1$ and $a_{1},\dots,a_{n}\in T$ such that
$f^{N}=a_{1}f_{1}+\cdots+a_{n}f_{n}.$ (2.1)
Let $s_{i}\in T_{f_{i}}$. By multiplying numerator and denominator, we may write each $s_{i}$ in the form $s_{i}=\frac{b_{i}}{f_{i}^{M}}$ for a common exponent $M$ (take $M$ larger than all individual exponents).
Step 2: candidate global section. Use the canonical maps $T_{f_{i}}\to T_{f}$ induced by the inclusions $D_{\Gamma}(f_{i})\subseteq D_{\Gamma}(f)$ and define
$s:=\sum_{i=1}^{n}a_{i}\cdot f_{i}\cdot s_{i}\cdot f^{-N}\ \in\ T_{f},$
where $f^{-N}$ denotes $\frac{1}{f^{N}}\in T_{f}$ and the sum is taken in the semiring $T_{f}$.
Step 3: verifying the restrictions. Fix $k\in\{1,\dots,n\}$. Restrict to $D_{\Gamma}(f_{k})$, i.e. map $T_{f}\to T_{f_{k}}$. In $T_{f_{k}}$, $f$ is invertible (since $D_{\Gamma}(f_{k})\subseteq D_{\Gamma}(f)$), so we may multiply by $f^{N}$. Using (2.1) we have $1=\sum_{i}a_{i}f_{i}f^{-N}$ in $T_{f_{k}}$. Therefore
$s|_{D_{\Gamma}(f_{k})}=\sum_{i}a_{i}f_{i}\big(s_{i}|_{D_{\Gamma}(f_{k})}\big)f^{-N}.$
On the overlap $D_{\Gamma}(f_{i})\cap D_{\Gamma}(f_{k})=D_{\Gamma}(f_{i}f_{k})$ we are given that $s_{i}$ and $s_{k}$ agree, so their images in $T_{f_{i}f_{k}}$ coincide. Since $T_{f_{i}f_{k}}$ receives maps from both $T_{f_{i}}$ and $T_{f_{k}}$, this implies that the image of $s_{i}$ in $T_{f_{k}}$ coincides with $s_{k}$ after localizing at $f_{i}$. In particular, the term $a_{i}f_{i}(s_{i}|_{D_{\Gamma}(f_{k})})$ equals $a_{i}f_{i}(s_{k})$ in $T_{f_{k}}$. Hence
$s|_{D_{\Gamma}(f_{k})}=\sum_{i}a_{i}f_{i}s_{k}f^{-N}=\Big(\sum_{i}a_{i}f_{i}f^{-N}\Big)s_{k}=1\cdot s_{k}=s_{k}.$
This proves existence.
Step 4: uniqueness. Suppose $s,t\in T_{f}$ restrict to the same section on every $D_{\Gamma}(f_{i})$. Write $s=\frac{p}{f^{r}}$ and $t=\frac{q}{f^{r}}$ after clearing denominators (common exponent $r$). The restriction equality in $T_{f_{i}}$ means that there exists $m_{i}$ such that
$f_{i}^{m_{i}}\cdot p=f_{i}^{m_{i}}\cdot q\quad\text{in }T,$
because equality in localization is detected by multiplication by a power of the localized element. Let $m:=\max_{i}m_{i}$. Then $f_{i}^{m}p=f_{i}^{m}q$ for every $i$. Multiply the relation (2.1) by itself $m$ times to obtain a decomposition
$f^{Nm}=\sum_{i}c_{i}f_{i}^{m}$
for suitable $c_{i}\in T$ (expand using distributivity). Multiplying by $p$ gives $f^{Nm}p=\sum_{i}c_{i}f_{i}^{m}p=\sum_{i}c_{i}f_{i}^{m}q=f^{Nm}q$. Thus $\frac{p}{f^{r}}=\frac{q}{f^{r}}$ in $T_{f}$ (multiply both numerators by $f^{Nm}$), hence $s=t$.
Definition 2.31 (Structure sheaf).
Since the $D_{\Gamma}(f)$ form a basis stable under finite intersections (Lemma 2.17), Theorem 2.29 implies that the presheaf $\mathcal{O}_{T}$ on principal opens extends uniquely to a sheaf on $\operatorname{Spec_{\Gamma}}(T)$, still denoted $\mathcal{O}_{T}$, by the usual sheafification/gluing procedure [25, 26]. The ringed space $(\operatorname{Spec_{\Gamma}}(T),\mathcal{O}_{T})$ is the affine $\Gamma$-scheme associated to $T$.
3 Affine $\Gamma$-schemes and the affine anti-equivalence
3.1 Affine $\Gamma$-schemes and morphisms
Definition 3.1 (Affine $\Gamma$-scheme).
An affine $\Gamma$-scheme is a ringed space of the form $(\operatorname{Spec_{\Gamma}}(T),\mathcal{O}_{T})$ constructed in Definition 2.31 from a commutative ternary $\Gamma$-semiring $T$. A (not necessarily affine) $\Gamma$-scheme is a ringed space locally isomorphic to affine $\Gamma$-schemes.
Definition 3.2 (Morphisms).
A morphism of $\Gamma$-schemes $f:(X,\mathcal{O}_{X})\to(Y,\mathcal{O}_{Y})$ is a morphism of ringed spaces such that on every affine open $V\simeq\operatorname{Spec_{\Gamma}}(S)\subseteq Y$ the induced map on sections respects the ternary $\Gamma$-operations: for all $U\subseteq f^{-1}(V)$ and all $a,b,c\in\mathcal{O}_{Y}(V)$ and $\gamma\in\Gamma$,
$f^{\#}_{V}\big(\{a\,b\,c\}_{\gamma}\big)=\{f^{\#}_{V}(a)\,f^{\#}_{V}(b)\,f^{\#}_{V}(c)\}_{\gamma}.$
3.2 Global sections on an affine $\Gamma$-scheme
Theorem 3.3 (Global sections recover the semiring).
For any commutative ternary $\Gamma$-semiring $T$, the canonical map
$\eta_{T}:\ T\longrightarrow\Gamma(\operatorname{Spec_{\Gamma}}(T),\mathcal{O}_{T}),\qquad a\longmapsto\big(a/1\text{ on each }D_{\Gamma}(f)\big),$
is an isomorphism of ternary $\Gamma$-semirings.
Proof 3.4.
Injectivity: if $a\neq b$ in $T$, then their images differ on $D_{\Gamma}(1)=\operatorname{Spec_{\Gamma}}(T)$ since $\mathcal{O}_{T}(D_{\Gamma}(1))=T$.
Surjectivity: let $s\in\Gamma(\operatorname{Spec_{\Gamma}}(T),\mathcal{O}_{T})$. By definition, $s$ is determined by its value on the cover $\operatorname{Spec_{\Gamma}}(T)=D_{\Gamma}(1)$. But $\mathcal{O}_{T}(D_{\Gamma}(1))=T_{1}\cong T$, so $s$ is represented by some element of $T$. Compatibility with $\Gamma$-ternary operations holds because the structure sheaf on principal opens is defined by localization of the underlying semiring, and localization maps preserve multiplication and the units $u_{\gamma}$.
3.3 Affine anti-equivalence
Theorem 3.5 (Affine anti-equivalence).
Let $T,S$ be commutative ternary $\Gamma$-semirings. Then there is a natural bijection
$\operatorname{Hom}_{\Gamma\text{-}\mathrm{TSR}}(T,S)\ \cong\ \operatorname{Hom}_{\Gamma\text{-}\mathrm{Sch}}\big(\operatorname{Spec_{\Gamma}}(S),\operatorname{Spec_{\Gamma}}(T)\big),$
contravariantly functorial in both variables. Moreover, for each $T$ the unit of this equivalence is the canonical isomorphism of Theorem 3.3.
Proof 3.6.
From algebra to geometry. Let $\varphi:T\to S$ be a ternary $\Gamma$-semiring homomorphism. Define $\varphi^{\ast}:\operatorname{Spec_{\Gamma}}(S)\to\operatorname{Spec_{\Gamma}}(T)$ by
$\varphi^{\ast}(P):=\varphi^{-1}(P).$
Since inverse images of prime ideals are prime, $\varphi^{\ast}$ is well-defined. Continuity follows from $\big(\varphi^{\ast}\big)^{-1}\big(D_{\Gamma}(f)\big)=D_{\Gamma}(\varphi(f))$.
For the sheaf map, for each principal open $D_{\Gamma}(f)\subseteq\operatorname{Spec_{\Gamma}}(T)$ we have a semiring homomorphism $T_{f}\to S_{\varphi(f)}$ induced by $\varphi$ and Proposition 2.26. These maps are compatible with restrictions, hence define a morphism of sheaves $\varphi^{\#}:\mathcal{O}_{T}\to(\varphi^{\ast})_{\ast}\mathcal{O}_{S}$, and the ternary $\Gamma$-compatibility is inherited from $\varphi(u_{\gamma})=u_{\gamma}$.
From geometry to algebra. Conversely, let $g:(\operatorname{Spec_{\Gamma}}(S),\mathcal{O}_{S})\to(\operatorname{Spec_{\Gamma}}(T),\mathcal{O}_{T})$ be a morphism of affine $\Gamma$-schemes. Apply global sections to obtain
$\Gamma(\operatorname{Spec_{\Gamma}}(T),\mathcal{O}_{T})\ \xrightarrow{g^{\#}}\ \Gamma(\operatorname{Spec_{\Gamma}}(S),\mathcal{O}_{S}).$
Using Theorem 3.3 on both sides yields a homomorphism $\varphi_{g}:T\to S$ of ternary $\Gamma$-semirings.
Mutual inverses. Starting from $\varphi$, the induced morphism on global sections recovers $\varphi$ because sections on $D_{\Gamma}(1)$ coincide with $T$ and $S$. Starting from $g$, the induced map on spectra agrees with $P\mapsto\varphi_{g}^{-1}(P)$ because on stalks the morphism of locally ringed spaces is determined by localization and hence by the induced map on global sections. Therefore the two constructions are inverse and natural.
4 Phase II: triadic symmetry and Nambu-type geometry
4.1 Triadic bracket on the structure sheaf
Fix the $\Gamma$-semiring units $u_{\gamma}$ as in Definition 2.2.
Definition 4.1 (Triadic bracket on sections).
Let $X=\operatorname{Spec_{\Gamma}}(T)$. For any open $U\subseteq X$ and any $\gamma\in\Gamma$, define
$\{\cdot,\cdot,\cdot\}_{\gamma,U}:\mathcal{O}_{T}(U)\times\mathcal{O}_{T}(U)\times\mathcal{O}_{T}(U)\to\mathcal{O}_{T}(U),\qquad(s_{1},s_{2},s_{3})\mapsto s_{1}\cdot s_{2}\cdot s_{3}\cdot u_{\gamma},$
where multiplication is taken in the semiring $\mathcal{O}_{T}(U)$ and $u_{\gamma}$ denotes the (global) unit section.
Lemma 4.2 (Restriction compatibility).
If $V\subseteq U$ are opens and $s_{i}\in\mathcal{O}_{T}(U)$, then
$\big(\{s_{1},s_{2},s_{3}\}_{\gamma,U}\big)\big|_{V}=\{s_{1}|_{V},s_{2}|_{V},s_{3}|_{V}\}_{\gamma,V}.$
Proof 4.3.
Restriction maps of $\mathcal{O}_{T}$ are semiring homomorphisms, hence preserve products and the distinguished unit sections $u_{\gamma}$.
4.2 A Nambu/Filippov-type identity
To align with Nambu-type geometry [13, 14] in a semiring context, we adopt the following “idempotent” variant, common in tropical/algebraic idempotent settings.
Definition 4.4 (Idempotent hypothesis).
A commutative semiring $(T,+,\cdot)$ is idempotent if $x+x=x$ for all $x\in T$. Equivalently, the additive monoid is a join-semilattice.
Definition 4.5 (Filippov fundamental identity (idempotent form)).
Let $(A,+)$ be an idempotent commutative monoid and let $\cdot,\cdot,\cdot]:A^{3}\to A$ be a trilinear operation. We say $\cdot$$\cdot,\cdot$] satisfies the (idempotent) Filippov identity if for all $x_{1},x_{2},y_{1},y_{2},y_{3}\in A$,
$x_{1},x_{2},[y_{1},y_{2},y_{3}]]=[[x_{1},x_{2},y_{1}],y_{2},y_{3}]+[y_{1},[x_{1},x_{2},y_{2}],y_{3}]+[y_{1},y_{2},[x_{1},x_{2},y_{3}]],$
where the right-hand side uses the addition of $A$.
Theorem 4.6 (Nambu-type identity for the triadic bracket).
Assume $T$ is idempotent in the sense of Definition $4.$ Then for every open $U\subseteq\operatorname{Spec_{\Gamma}}(T)$ and every fixed $\gamma\in\Gamma$, the triadic bracket $\cdot,\cdot,\cdot]:=\{\cdot,\cdot,\cdot\}_{\gamma,U}$ on $\mathcal{O}_{T}(U)$ satisfies the idempotent Filippov identity (Definition 4.5).
Proof 4.7.
Let $\cdot,\cdot,\cdot]=\{\cdot,\cdot,\cdot\}_{\gamma,U}$. Because the bracket is defined by multiplication and $u_{\gamma}$ is central, each of the three terms on the right-hand side equals the same product
$x_{1}\cdot x_{2}\cdot y_{1}\cdot y_{2}\cdot y_{3}\cdot u_{\gamma}^{2}$
(up to rebracketing and commutativity). The left-hand side equals the same product. Since addition is idempotent, the sum of three identical terms equals that term.
4.3 Functoriality and invariance
Definition 4.8 ($\Gamma$-automorphisms).
Let $\operatorname{Aut}_{\Gamma}(T)$ denote the group of ternary $\Gamma$-semiring automorphisms of $T$ (i.e. semiring automorphisms preserving each unit $u_{\gamma}$).
Proposition 4.9 (Action on spectrum and sheaf).
Each $\sigma\in\operatorname{Aut}_{\Gamma}(T)$ induces:
  1. 1.
    a homeomorphism $\sigma^{\ast}:\operatorname{Spec_{\Gamma}}(T)\to\operatorname{Spec_{\Gamma}}(T)$, $P\mapsto\sigma^{-1}(P)$, and
  2. 2.
    an isomorphism of sheaves of ternary $\Gamma$-semirings $\sigma^{\#}:\mathcal{O}_{T}\to(\sigma^{\ast})_{\ast}\mathcal{O}_{T}$.
Moreover, $\sigma^{\#}$ preserves the triadic bracket of Definition 4.1.
Proof 4.10.
The map on primes is well-defined and continuous by the same argument as in Theorem 3.5. On principal opens, $\sigma$ induces compatible homomorphisms on localizations $T_{f}\to T_{\sigma(f)}$, hence a sheaf morphism. Bracket invariance holds because $\sigma$ preserves multiplication and $\sigma(u_{\gamma})=u_{\gamma}$.
Theorem 4.11 (Localization and stalk compatibility).
For each $P\in\operatorname{Spec_{\Gamma}}(T)$, the stalk $\mathcal{O}_{T,P}$ inherits a triadic bracket, and the canonical localization map $T\to\mathcal{O}_{T,P}$ is a homomorphism of ternary $\Gamma$-semirings.
Proof 4.12.
The stalk is a filtered colimit of the semirings $T_{f}$ over all $f\notin P$. Each $T_{f}$ carries the induced bracket by Definition 4.1 on $D_{\Gamma}(f)$. Compatibility with restriction (Lemma 4.2) implies these brackets are compatible in the colimit, hence descend to a bracket on the stalk.
5 Phase III: canonical Laplacian and spectral decomposition
5.1 Finite spectra and the specialization graph
In full generality, $\operatorname{Spec_{\Gamma}}(T)$ can be infinite. The Laplacian construction below is canonical on finite spectra; this is the setting in which explicit eigenvalue computations and spectral factorization are meaningful.
Definition 5.1 (Specialization order).
Let $X=\operatorname{Spec_{\Gamma}}(T)$. For $P,Q\in X$, write $P\leq Q$ if $P\subseteq Q$ (equivalently, $Q\in\operatorname{cl}(\{P\})$). This is the specialization preorder, and for $X$ a $T_{0}$ space it is a partial order.
Definition 5.2 (Specialization graph and Laplacian).
Assume $X=\operatorname{Spec_{\Gamma}}(T)$ is finite. Define an undirected graph $G_{X}$ with vertex set $X$ and an edge between distinct vertices $P\neq Q$ iff $P$ and $Q$ are comparable in the specialization order (i.e. $P\subseteq Q$ or $Q\subseteq P$). Let $A_{X}$ be its adjacency matrix and $D_{X}$ the diagonal matrix of degrees. The canonical Laplacian is
$L_{X}:=D_{X}-A_{X}.$
Remark 5.3.
The construction depends only on the underlying finite $T_{0}$ space $(X,\leq)$. It is therefore invariant under isomorphisms of affine $\Gamma$-schemes and is well-suited to “geometric connected components” detectable from specialization structure.
5.2 Isomorphism invariance
Theorem 5.4 (Invariance under affine $\Gamma$-scheme isomorphisms).
Let $\psi:(\operatorname{Spec_{\Gamma}}(S),\mathcal{O}_{S})\to(\operatorname{Spec_{\Gamma}}(T),\mathcal{O}_{T})$ be an isomorphism of affine $\Gamma$-schemes and assume both spectra are finite. Then the induced bijection on points is an isomorphism of specialization graphs, hence
$L_{\operatorname{Spec_{\Gamma}}(S)}=P^{-1}L_{\operatorname{Spec_{\Gamma}}(T)}P$
for a permutation matrix $P$ corresponding to the vertex relabeling.
Proof 5.5.
A scheme isomorphism is, in particular, a homeomorphism. For spectral spaces, homeomorphisms preserve the specialization order, hence preserve comparability. Therefore the adjacency relation is preserved and the Laplacians are permutation-similar.
5.3 Block-diagonal decomposition under clopen components
Theorem 5.6 (Block decomposition from clopen components).
Let $X=\operatorname{Spec_{\Gamma}}(T)$ be finite and suppose $X$ decomposes as a disjoint union of clopen subsets
$X=X_{1}\sqcup\cdots\sqcup X_{r}.$
Then the specialization graph $G_{X}$ is the disjoint union of the induced subgraphs $G_{X_{i}}$, and the Laplacian is permutation-similar to a block diagonal matrix
$L_{X}\sim\begin{pmatrix}L_{X_{1}}&0&\cdots&0\\ 0&L_{X_{2}}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&L_{X_{r}}\end{pmatrix}$.
Proof 5.7.
If $X_{i}$ is clopen, then it is closed under specialization and generalization: if $P\in X_{i}$ and $P\leq Q$ or $Q\leq P$, then $Q\in\operatorname{cl}(\{P\})\subseteq X_{i}$ because $X_{i}$ is closed, and similarly for open. Thus no comparability relation can occur between points in different $X_{i}$’s. Hence $G_{X}$ has no edges across components and is the disjoint union of the induced subgraphs. The Laplacian of a disjoint union is block diagonal up to relabeling of vertices.
5.4 Algebraic connectivity and geometric connectedness
We now link the second Laplacian eigenvalue (Fiedler value) to connectedness. The graph-theoretic statement is classical [15, 16].
Theorem 5.8 (Fiedler connectivity criterion).
Let $G$ be a finite undirected graph with Laplacian $L$. Let $0=\lambda_{1}\leq\lambda_{2}\leq\cdots\leq\lambda_{n}$ be the eigenvalues of $L$. Then $\lambda_{2}>0$ if and only if $G$ is connected.
Theorem 5.9 (Connectedness of finite $\Gamma$-spectra via $\lambda_{2}$).
Let $X=\operatorname{Spec_{\Gamma}}(T)$ be finite. Then $X$ is connected (as a topological space) if and only if $\lambda_{2}(L_{X})>0$.
Proof 5.10.
For a finite $T_{0}$ space, connectedness is equivalent to connectedness of the undirected graph obtained from the specialization order (equivalently, the Hasse diagram viewed undirected). Indeed, if $X$ decomposes into two disjoint clopen subsets, there is no specialization relation between them, so the comparability graph disconnects; conversely, if the comparability graph disconnects, the vertex set splits into two subsets closed under specialization and generalization, which are therefore clopen.
Apply Theorem 5.8 to the specialization graph $G_{X}$.
6 Phase IV: explicit computed finite examples
All examples below are finite idempotent semirings (hence Theorem 4.6 applies), and we take $\Gamma$ to be the trivial group so that $u_{\gamma}=1$. The ternary operation is $\{abc\}:=abc$ (binary product of three elements), which in these idempotent examples is the meet $\wedge$.
6.1 Example 1: a connected two-point spectrum (Sierpiński space)
Example 6.1 (Three-element chain semiring).
Let $T=\{0,e,1\}$ with addition $x+y:=\max\{x,y\}$ (join) and multiplication $x\cdot y:=\min\{x,y\}$ (meet), with order $0<e<1$. This is a commutative idempotent semiring (a distributive lattice).
Proposition 6.2.
The prime ideals of $T$ are $P_{0}=\{0\}$ and $P_{1}=\{0,e\}$. Thus $\operatorname{Spec_{\Gamma}}(T)=\{P_{0},P_{1}\}$ with specialization $P_{0}\subset P_{1}$, and the topology is the Sierpiński topology.
Proof 6.3.
Ideals in a finite chain are down-sets. The proper ideals are $\{0\}$ and $\{0,e\}$. Both are prime: if $x\wedge y\in\{0\}$ then $x=0$ or $y=0$; if $x\wedge y\in\{0,e\}$ then $x\leq e$ or $y\leq e$.
Proposition 6.4 (Canonical Laplacian).
The specialization graph is a single edge $P_{0}$$P_{1}$. Hence
$A=\begin{pmatrix}0&1\\ 1&0\end{pmatrix},\qquad D=\begin{pmatrix}1&0\\ 0&1\end{pmatrix},\qquad L=D-A=\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$.
The eigenvalues are $0,2$ with eigenvectors $(1,1)$ and (1,1)(1,-1).
Proof 6.5.
There is an edge because $P_{0}\subset P_{1}$. The eigen-computation is immediate.
6.2 Example 2: a disconnected spectrum and exact block decomposition
Example 6.6 (Product of Boolean semirings).
Let $B=\{0,1\}$ with ++ as $\vee$ and $\cdot$ as $\wedge$. Set $T=B\times B$ with componentwise operations.
Proposition 6.7.
The prime ideals of $T$ are
$P_{1}=\{0\}\times B,\qquad P_{2}=B\times\{0\}.$
The $\Gamma$-spectrum is the discrete two-point space $\{P_{1},P_{2}\}$, hence disconnected.
Proof 6.8.
Prime ideals of a finite product correspond to primes in each factor. In $B$ the unique proper prime ideal is $\{0\}$. Taking inverse images under the projections yields $P_{1}$ and $P_{2}$. Both are maximal and incomparable, so the topology is discrete.
Proposition 6.9 (Laplacian and spectrum).
The specialization graph has no edges (the two primes are incomparable), hence
$L=\begin{pmatrix}0&0\\ 0&0\end{pmatrix},\qquad\text{eigenvalues }0,0.$
This is the block decomposition of Theorem 5.6 with two clopen components.
Proof 6.10.
No comparability means no adjacency, hence $A=0$ and $L=0$.
6.3 Example 3: a three-point connected spectrum with nontrivial algebraic connectivity
Example 6.11 (Four-element chain semiring).
Let $T=\{0<a<b<1\}$ with join as addition and meet as multiplication (a four-element chain lattice).
Proposition 6.12.
The prime ideals are
$P_{0}=\{0\},\qquad P_{1}=\{0,a\},\qquad P_{2}=\{0,a,b\},$
forming a chain $P_{0}\subset P_{1}\subset P_{2}$. Thus $\operatorname{Spec_{\Gamma}}(T)$ is connected and its specialization graph is the complete graph $K_{3}$.
Proof 6.13.
As in Proposition 6.2, the proper ideals are down-sets. Each of the three displayed ideals is prime by the chain argument. Any two of them are comparable, hence the comparability graph is $K_{3}$.
Proposition 6.14 (Laplacian and eigenvalues).
With vertex order $(P_{0},P_{1},P_{2})$, the adjacency and Laplacian matrices are
$A=\begin{pmatrix}0&1&1\\ 1&0&1\\ 1&1&0\end{pmatrix},\qquad D=\begin{pmatrix}2&0&0\\ 0&2&0\\ 0&0&2\end{pmatrix},\qquad L=D-A=\begin{pmatrix}2&-1&-1\\ -1&2&-1\\ -1&-1&2\end{pmatrix}$.
The eigenvalues are $0,3,3$. In particular, $\lambda_{2}=3>0$, confirming connectedness by Theorem 5.9.
Proof 6.15.
The Laplacian of $K_{3}$ is well known [16]. A direct computation gives eigenvalues $0,3,3$.
Appendix A Optional application: fuzzy robustness viewpoint
This appendix is intentionally conservative: it records only basic definitions and a stability statement needed to interpret “robustness” of prime separation under small perturbations, without impacting the main SIGMA narrative.
Definition A.1 (Fuzzy subset and $\alpha$-cuts).
A fuzzy subset of $T$ is a function $\mu:T\to[0,1]$ [20]. For $\alpha\in(0,1]$, its $\alpha$-cut is the crisp subset
$\mu]_{\alpha}:=\{x\in T\mid\mu(x)\geq\alpha\}.$
Definition A.2 (Fuzzy $\Gamma$-ideal (one standard choice)).
A fuzzy subset $\mu:T\to[0,1]$ is a fuzzy $\Gamma$-ideal if for all $x,y,z\in T$ and $\gamma\in\Gamma$:
  1. 1.
    $\mu(0)=1$ and $\mu(x+y)\geq\min\{\mu(x),\mu(y)\}$,
  2. 2.
    $\mu(\{x\,y\,z\}_{\gamma})\geq\mu(x)$.
Remark A.3.
Many equivalent or stronger definitions exist in the fuzzy algebra literature [22, 23, 24]. The above is sufficient for the elementary stability lemma below.
Lemma A.4 (Sup-norm stability of $\alpha$-cuts).
Let $\mu,\nu:T\to[0,1]$ be fuzzy subsets and set $\|\mu-\nu\|_{\infty}:=\sup_{x\in T}|\mu(x)-\nu(x)|$. Then for every $\alpha\in(0,1]$ and every $\varepsilon>0$,
$\|\mu-\nu\|_{\infty}\leq\varepsilon\quad\Longrightarrow\quad[\mu]_{\alpha+\varepsilon}\subseteq[\nu]_{\alpha}\subseteq[\mu]_{\alpha-\varepsilon},$
where $\mu]_{\beta}:=\emptyset$ if $\beta>1$ and $\mu]_{\beta}:=T$ if $\beta\leq 0$.
Proof A.5.
If $\mu(x)\geq\alpha+\varepsilon$ and $|\mu(x)-\nu(x)|\leq\varepsilon$, then $\nu(x)\geq\alpha$. The other inclusion is analogous.
Appendix B Optional appendix: a conservative spectral clustering pipeline
Given a finite $\Gamma$-spectrum $X$ and its canonical Laplacian $L_{X}$ (Definition 5.2), one may apply standard spectral clustering heuristics [18, 19]. We include this only as an illustrative computational layer (not a mathematical claim about learning).
  1. 1.
    Compute the eigenpairs of $L_{X}$ (exactly in small examples; numerically in general) [17].
  2. 2.
    Choose $k$ and form the matrix $U\in\mathbb{R}^{|X|\times k}$ whose columns are eigenvectors corresponding to the $k$ smallest eigenvalues.
  3. 3.
    Normalize the rows of $U$ and run $k$-means on the embedded points.
  4. 4.
    Interpret clusters as a coarse partition of $X$; if $X$ is disconnected, Theorem 5.6 yields an exact decomposition already at the level of $L_{X}$.

### v2

@@
 
-[…]
 […]
@@
     for all $a\in I$, all $b,c\in T$, and all $\gamma\in\Gamma$, one has $\{a\,b\,c\}_{\gamma}\in I$.
-Equivalently, $I$ is an ideal of the underlying semiring $(T,+,\cdot,0,1)$ (since $u_{\gamma}$ is a unit for every $\gamma$).
+Equivalently, II is an ideal of the underlying semiring $(T,+,\cdot,0,1)$ (since $u_{\gamma}$ is a unit for every $\gamma$).
 Definition 2.5 (Generated $\Gamma$-ideal).
@@
 Definition 2.6 ($\Gamma$-radical).
-For a $\Gamma$-ideal $I$, define its $\Gamma$-radical by
+For a $\Gamma$-ideal II, define its $\Gamma$-radical by
 $\sqrt[\Gamma]{\,\cdot\,}(I):=\{x\in T\mid\exists n\geq 1\text{ with }x^{n}\in I\},$
@@
 Lemma 2.7.
-For any $\Gamma$-ideal $I$, the set $\sqrt[\Gamma]{\,\cdot\,}(I)$ is a $\Gamma$-ideal containing $I$.
+For any $\Gamma$-ideal II, the set $\sqrt[\Gamma]{\,\cdot\,}(I)$ is a $\Gamma$-ideal containing II.
 Proof 2.8.
 If $x^{n}\in I$ then clearly $x\in\sqrt[\Gamma]{\,\cdot\,}(I)$; hence $I\subseteq\sqrt[\Gamma]{\,\cdot\,}(I)$.
-Closure under addition and multiplication by arbitrary elements is standard: if $x^{m}\in I$ and $y^{n}\in I$, then $(x+y)^{m+n}$ expands into a sum of monomials each divisible by $x^{m}$ or $y^{n}$, hence lies in $I$; similarly $(t\cdot x)^{m}=t^{m}x^{m}\in I$.
+Closure under addition and multiplication by arbitrary elements is standard: if $x^{m}\in I$ and $y^{n}\in I$, then $(x+y)^{m+n}$ expands into a sum of monomials each divisible by $x^{m}$ or $y^{n}$, hence lies in II; similarly $(t\cdot x)^{m}=t^{m}x^{m}\in I$.
 Since $\Gamma$-ideal closure is equivalent to ideal closure in our setting (Definition 2.4), we are done.
@@
 Lemma 2.11 (Radical as intersection of primes).
-Let $I$ be a $\Gamma$-ideal. Then
+Let II be a $\Gamma$-ideal. Then
 $\sqrt[\Gamma]{\,\cdot\,}(I)=\bigcap_{\begin{subarray}{c}P\ \text{prime }\Gamma\text{-ideal}\\ I\subseteq P\end{subarray}}P.$
 Proof 2.12.
-($\subseteq$) If $x\in\sqrt[\Gamma]{\,\cdot\,}(I)$, pick $n$ with $x^{n}\in I$. If $P$ is prime and contains $I$, then $x^{n}\in P$ implies $x\in P$. So $x$ lies in every prime containing $I$.
+($\subseteq$) If $x\in\sqrt[\Gamma]{\,\cdot\,}(I)$, pick $n$ with $x^{n}\in I$. If $P$ is prime and contains II, then $x^{n}\in P$ implies $x\in P$. So $x$ lies in every prime containing II.
 ($\supseteq$) Suppose $x\notin\sqrt[\Gamma]{\,\cdot\,}(I)$. Consider the collection
@@
 Hence $a\in P$ or $b\in P$, proving primeness.
-Finally, $I\subseteq P$ by construction, and $x\notin P$ because $x\in P$ would force $x^{1}\in P$, contradicting $P\in\mathcal{S}$. Thus $x$ is not in the intersection of primes containing $I$.
+Finally, $I\subseteq P$ by construction, and $x\notin P$ because $x\in P$ would force $x^{1}\in P$, contradicting $P\in\mathcal{S}$. Thus $x$ is not in the intersection of primes containing II.
 2.3 The $\Gamma$-spectrum and $\Gamma$-Zariski topology
@@
 Definition 2.14 (Closed sets and principal opens).
-For a $\Gamma$-ideal $I$, define
+For a $\Gamma$-ideal II, define
 $V(I):=\{P\in\operatorname{Spec_{\Gamma}}(T)\mid I\subseteq P\}.$
@@
 ($\Rightarrow$) Let $I=\langle f_{1},\dots,f_{n}\rangle_{\Gamma}$.
-The equality of opens says: for every prime $P$, if $f\notin P$ then $f_{i}\notin P$ for some $i$.
+The equality of opens says: for every prime $P$, if $f\notin P$ then $f_{i}\notin P$ for some ii.
 Equivalently, if $f_{i}\in P$ for all $i$ (i.e.
 $I\subseteq P$), then $f\in P$.
-Thus $f$ lies in the intersection of all primes containing $I$, hence $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ by Lemma 2.11.
+Thus $f$ lies in the intersection of all primes containing II, hence $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ by Lemma 2.11.
 ($\Leftarrow$) Suppose $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$.
 Let $P\in D_{\Gamma}(f)$, so $f\notin P$.
-If $P$ contained $I$, then $f\in\sqrt[\Gamma]{\,\cdot\,}(I)\subseteq\sqrt[\Gamma]{\,\cdot\,}(P)=P$, contradicting $f\notin P$.
+If $P$ contained II, then $f\in\sqrt[\Gamma]{\,\cdot\,}(I)\subseteq\sqrt[\Gamma]{\,\cdot\,}(P)=P$, contradicting $f\notin P$.
 Hence $I\not\subseteq P$, so some generator $f_{i}\notin P$, i.e.
 $P\in D_{\Gamma}(f_{i})$.
-Thus $D_{\Gamma}(f)\subseteq\bigcup_{i}D_{\Gamma}(f_{i})$, and the reverse inclusion holds because each $D_{\Gamma}(f_{i})\subseteq D_{\Gamma}(f)$ is automatic from the displayed radical inclusion: $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ implies $D_{\Gamma}(f)\supseteq D_{\Gamma}(f_{i})$ for each $i$.
+Thus $D_{\Gamma}(f)\subseteq\bigcup_{i}D_{\Gamma}(f_{i})$, and the reverse inclusion holds because each $D_{\Gamma}(f_{i})\subseteq D_{\Gamma}(f)$ is automatic from the displayed radical inclusion: $f\in\sqrt[\Gamma]{\,\cdot\,}(I)$ implies $D_{\Gamma}(f)\supseteq D_{\Gamma}(f_{i})$ for each ii.
 Corollary 2.21.
@@
 Proof 2.22.
-From Theorem 2.19 we have $f\in\sqrt[\Gamma]{\,\cdot\,}(\langle f_{1},\dots,f_{n}\rangle_{\Gamma})$, so $f^{N}\in\langle f_{1},\dots,f_{n}\rangle_{\Gamma}$ for some $N$.
+From Theorem 2.19 we have $f\in\sqrt[\Gamma]{\,\cdot\,}(\langle f_{1},\dots,f_{n}\rangle_{\Gamma})$, so $f^{N}\in\langle f_{1},\dots,f_{n}\rangle_{\Gamma}$ for some NN.
 By definition of generated ideal, $f^{N}$ is a finite sum of multiples of the generators, giving the claim.
@@
   2. 1.
-    (existence of gluing) For any sections $s_{i}\in T_{f_{i}}$ such that their restrictions agree on each overlap $D_{\Gamma}(f_{i})\cap D_{\Gamma}(f_{j})=D_{\Gamma}(f_{i}f_{j})$, there exists $s\in T_{f}$ with $s|_{D_{\Gamma}(f_{i})}=s_{i}$ for all $i$.
+    (existence of gluing) For any sections $s_{i}\in T_{f_{i}}$ such that their restrictions agree on each overlap $D_{\Gamma}(f_{i})\cap D_{\Gamma}(f_{j})=D_{\Gamma}(f_{i}f_{j})$, there exists $s\in T_{f}$ with $s|_{D_{\Gamma}(f_{i})}=s_{i}$ for all ii.
   3. 2.
@@
 Let $m:=\max_{i}m_{i}$.
-Then $f_{i}^{m}p=f_{i}^{m}q$ for every $i$.
+Then $f_{i}^{m}p=f_{i}^{m}q$ for every ii.
 Multiply the relation (2.1) by itself $m$ times to obtain a decomposition
@@
 $\operatorname{Hom}_{\Gamma\text{-}\mathrm{TSR}}(T,S)\ \cong\ \operatorname{Hom}_{\Gamma\text{-}\mathrm{Sch}}\big(\operatorname{Spec_{\Gamma}}(S),\operatorname{Spec_{\Gamma}}(T)\big),$
-contravariantly functorial in both variables. Moreover, for each $T$ the unit of this equivalence is the canonical isomorphism of Theorem 3.3.
+contravariantly functorial in both variables. Moreover, for each $T$ the unit of this equivalence is the canonical isomorphism of Theorem $3.$
 Proof 3.6.
@@
 4.1 Triadic bracket on the structure sheaf
-Fix the $\Gamma$-semiring units $u_{\gamma}$ as in Definition 2.2.
+Fix the $\Gamma$-semiring units $u_{\gamma}$ as in Definition $2.$
 Definition 4.1 (Triadic bracket on sections).
@@
 $x_{1},x_{2},[y_{1},y_{2},y_{3}]]=[[x_{1},x_{2},y_{1}],y_{2},y_{3}]+[y_{1},[x_{1},x_{2},y_{2}],y_{3}]+[y_{1},y_{2},[x_{1},x_{2},y_{3}]],$
-where the right-hand side uses the addition of $A$.
+where the right-hand side uses the addition of AA.
 Theorem 4.6 (Nambu-type identity for the triadic bracket).
@@
 for a permutation matrix $P$ corresponding to the vertex relabeling.
-Proof 5.5.
+Proof $5.$
 A scheme isomorphism is, in particular, a homeomorphism.